# Sparse High Rank Adapters

### v1

@@ Abstract
Low Rank Adaptation (LoRA) has gained massive attention in the recent generative AI research. One of the main advantages of LoRA is its ability to be fused with pretrained models adding no overhead during inference. However, from a mobile deployment standpoint, we can either avoid inference overhead in the fused mode but lose the ability to switch adapters rapidly, or suffer significant (up to 30% higher) inference latency while enabling rapid switching in the unfused mode. LoRA also exhibits concept-loss when multiple adapters are used concurrently. In this paper, we propose Sparse High Rank Adapters (SHiRA), a new paradigm which incurs no inference overhead, enables rapid switching, and significantly reduces concept-loss. Specifically, SHiRA can be trained by directly tuning only $1$-$2\%$ of the base model weights while leaving others unchanged. This results in a highly sparse adapter which can be switched directly in the fused mode. We further provide theoretical and empirical insights on how high sparsity in SHiRA can aid multi-adapter fusion by reducing concept loss. Our extensive experiments on LVMs and LLMs demonstrate that finetuning only a small fraction of the parameters in the base model is sufficient for many tasks while enabling both rapid switching and multi-adapter fusion. Finally, we provide a latency- and memory-efficient SHiRA implementation based on Parameter-Efficient Finetuning (PEFT) Library. This implementation trains at nearly the same speed as LoRA while consuming lower peak GPU memory, thus making SHiRA easy to adopt for practical use cases.
@@ 1 Introduction
Low Rank Adaptation (LoRA) [11] is an established technique to tune the behavior of large generative models such as Large Language Models (LLMs) [28, 27] and Stable Diffusion [22, 20]. As the name suggests, LoRA requires very few parameters since it trains low rank projection weights that consume very low memory during the finetuning process while producing excellent results. Moreover, these low rank weights can be fused analytically into the base model weights, thereby incurring no additional overhead during inference.
@@
Figure 1: Sparse High Rank Adapters (SHiRA): Changing about $1$-$2\%$ weights of the pretrained generative model is often sufficient to achieve high performance. Due to its extreme sparsity, SHiRA enables rapid switching and also reduced concept loss during multi-adapter fusion. In contrast, LoRA modifies majority of parameters when fused, prohibiting rapid switching on mobile devices and also experiences concept loss during multi-adapter fusion. For LoRA, elephant for single “paintings” adapter case has artifacts (extra/broken tusks), and bird/knight for multi-adapter case lose “paintings” concept and keep only the “blue fire” effects. SHiRA does not experience these issues.
@@
  • We propose SHiRA, a new high rank adapter paradigm and demonstrate that changing as few as $1$-$2\%$ parameters of the original network is sufficient for many adaptation tasks.
@@
  • Finally, on the training side, we provide a PEFT-based latency- and memory-efficient implementation for SHiRA which trains nearly as fast as standard LoRA while consuming $16\%$ lower peak GPU memory. Beyond PEFT, we provide a simple way to turn any trainer into SHiRA finetuning. For inference, we demonstrate that SHiRA weights can be loaded on a CPU up to 10×10$\times$10 × faster than equivalent LoRA fusing, thereby enabling rapid switching.
The rest of this paper is organized as follows: section 2 presents the background and related work. We propose SHiRA in section 3 while describing its theoretical properties in section 4. We then conduct extensive experiments for SHiRA in section 5. Finally, we conclude the paper in section 6.
@@ 2.2 Related Work
Our work is most related to the highly popular LoRA, its variants, sparse adapters, partial finetuning, and multi-adapter fusion techniques, as discussed below.
@@
Partial Finetuning. Our work is most closely related to partial finetuning techniques that were mostly proposed in the pre-LoRA era [34, 26, 3, 31, 8]. These methods use a mix of fixed sparse masks [26] or learned masks [34, 8] to finetune a pretrained network. Note that, these techniques have been mostly explored for relatively small language models, and not for recent LLMs and diffusion models. Since the LoRA models exploded in popularity, it has been unclear if other sparse finetuning techniques would achieve comparable results to LoRA on generic adapter tasks, particularly in the vision domain. One significant limitation of partial finetuning, as opposed to LoRA-based methods, is its high GPU memory consumption, making it impractical to be used for large generative models. Consequently, the reduced memory consumption for finetuning was a key factor to LoRA’s success and its widespread adoption. To this end, we provide a memory- and latency-efficient implementation for SHiRA which trains as efficiently as LoRA, thus requiring significantly lower memory consumption compared to prior partial finetuning techniques. Further, we explore the effectiveness of sparse finetuning on both large language and vision models and provide a detailed analysis on rapid switching and multi-adapter fusion of the high rank adapters. To our knowledge, high rank characteristics of sparse adapters has not been addressed by the partial finetuning literature.
@@
Figure 2: (a) LoRA appends two low rank weights that can be fused into the pretrained weights at inference time. However, this modifies all weights and prevents rapid adapter switching. (b) SHiRA does not require additional weights during training but finetunes very few pretrained weights. Our approach relies on a sparse mask for gradient-masking during training. We show that finetuning as low as $1$-$2\%$ parameters is sufficient to achieve high accuracy on many adapter tasks.
@@ 3.1 Sparse High Rank Adapters (SHiRA)
SHiRA exploits highly sparse trainable parameters in the pretrained base model. In its simplest form, our proposed adapter can be trained by masking gradients such that only a fraction of original weights get updated. Specifically, we do not add any new weights to the forward pass like LoRA (see Fig. 2(a)) but rather make a small percentage of existing weights trainable (see Fig. 2(b) top). To this end, we first create an extremely sparse ($\sim$98989898-$99\%$ zeros) mask $\mathcal{M}\in\mathbb{R}^{n\times m}=\{0,1\}^{n\times m}$, where $n,m$ are dimensions of the pretrained weight matrix. $\mathcal{M}$ is then used to mask the gradients during backpropagation using a Hadamard product (see Fig. 2(b) bottom). As a result, very few parameters get updated during training and our adapter consists of just those sparse weights. We will provide concrete gradient masking-based and another latency-/memory-efficient PEFT-based implementations for SHiRA in section 3.3.
Various strategies to create the mask $\mathcal{M}$ are described below (more details are in Appendix B). Note that, we only set $1$-$2\%$ of the parameters as trainable in each case:
@@
    SHiRA-SNIP: Finally, this mask is based on SNIP metric used in the pruning literature [17]. Here we combine weight magnitude and gradient strategies, i.e., SNIP equals gradient magnitude times the weight magnitude.
@@ 3.2 Rapid Adapter Switching, Multi-Adapter Fusion, and High Rank
Since very few base weights change during the SHiRA training, we can simply extract them out and store them as sparse weights and their indices (see Fig. 3(a)). Hence, SHiRA is comparable to LoRA in model size but overwrites only a fraction of the pretrained weights at inference time. In contrast, LoRA fuses into base weights as $W_{new}=W+AB$ and changes the entire weight. Note that, we do not actually need to fuse SHiRA but rather just need to overwrite the modified value at the correct index in the pretrained weight tensor. This enables rapid switching on resource-constrained devices. To verify that SHiRA indeed provides rapid switching benefits compared to LoRA, we provide an optimized implementation based on scatter_op to overwrite base model weights instead of fusing them like LoRA. We demonstrate that on a CPU, weight loading for SHiRA adapters can be up to $\bm{10\times}$ faster than equivalent LoRA fusing for inference (see Appendix C and Fig 7).
Next, we discuss multi-adapter fusion in SHiRA. Suppose, we are given two adapters $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ with sparse masks $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. There are two important aspects to consider here: (i) What is the impact of sparsity on relative interference between adapters in the multi-adapter setting? (ii) Is it possible to create certain types of masks that result in nearly orthogonal SHiRA weights so they do not significantly interfere with each other at inference time?
@@
Finally, since we do not have any low rank weights in the forward pass, our proposed adapters can be high rank albeit highly sparse. We show that this high sparsity with high rank has interesting theoretical properties in the next section.
@@ 4 Theoretical Insights for SHiRA
4.1 Sparse High Rank Adapters
@@ 4.2 Adapter Weight Orthogonality in Multi-Adapter Fusion
Non-overlapping SHiRA-Struct adapters are nearly orthogonal. That is, AWOR for non-overlapping SHiRA-Struct adapters is at most the sum of sparsity of individual adapters. Since all SHiRA masks are highly sparse, this means that the product $\mathcal{A}_{1}^{T}\mathcal{A}_{2}$ has a lot of zeros, thus making the adapters nearly orthogonal.
@@
Remark 1. The orthogonality property shown here can lead to disentangled representation for adapter outputs before they merge into the base model. However, this property does not hold for other SHiRA masks that do not have a regular sparsity pattern like SHiRA-Struct even if other SHiRA strategies are still more orthogonal than LoRA weights (e.g., see SHiRA-WM AWOR in Fig. 4(right)). Importantly, this analysis only focuses on orthogonality of adapter weights and not on orthogonality of subspaces. We leave the subspace analysis of SHiRA for future work.
@@ 5 Experiments
In this section, we showcase the effectiveness of SHiRA and conduct extensive evaluation on various tasks and benchmarks in both vision and language domains.
@@ 5.2.1 Impact of Various SHiRA Masks
Next, we compare various types of SHiRA masks in Fig. 5. Clearly, all SHiRA schemes produce impressive images for different prompts and significantly outperform LoRA. We further quantify the image quality using HPSv2 for each of the masks. The results are presented in Table 1. As evident, all variants of SHiRA consistently achieve superior or similar HPSv2 scores than LoRA, especially for larger $\alpha$ (see details on scaling factor $\alpha$ in Appendix I). More results are provided in Appendices J and K: see Table 9 and Fig.  9,  1011.
@@ 5.2.2 SHiRA Adapters aid Multi-Adapter Fusion
We now validate the effectiveness of various SHiRA schemes on multi-adapter fusion. The right two columns in Fig. 1 and Fig. 5 show our results. SHiRA is clearly better at capturing both concepts than LoRA. For example, both bird and knight images in Fig. 1 generated with LoRA lose most of the paintings concept. Similarly, the fox image in Fig. 5 where LoRA does not show significant bluefire concept. In contrast, SHiRA-Struct and SHiRA-SNIP consistently perform well on many different prompts and produce exceptional images for multi-adapter fusion. Please refer to Appendix K (Fig. 91011, and  12) for additional results. For certain classes that were not included in the training set for both adapters (e.g., see Koala in Fig. 911, and 12), we observe that LoRA produces significant artifacts whereas SHiRA generates high quality images.
@@ 5.3.1 Single Adapter SHiRA Finetuning
Similar to vision results, we demonstrate the effectiveness of SHiRA on language tasks. For our experiments, each adapter (i.e., weight-magnitude, gradient-magnitude and SNIP based SHiRA) is trained on the combined 170K sample commonsense reasoning dataset released by [12, 18]. Similar to [18], we train our SHiRA adapters for 3 epochs and compare it against the LoRA baselines. As shown in Table 2, various SHiRA adapters outperform LoRA by 1.9-2.7% on an average on LLaMA-7B. Importantly, SHiRA only modifies 1% base parameter weights as compared to 66.25% (4.5B weights) changed by LoRA in the fused mode, thus enabling rapid switching on edge devices. Interestingly, we found that SHiRA-Struct does not perform well on language tasks likely because it is a rank 1 + diagonal adapter and may not have sufficient expressive power.
@@
Table 2: Evaluation of LLaMA-7B models on Commonsense Reasoning. WG and HS denote WinoGrande and HellaSwag, respectively. %C represents parameters changed in the fused mode. ($\uparrow$): the higher the better. Green denotes improvement over baselines. Trained by masking a high-rank DoRA with a WM mask of top 1% weights, thus changing only 1% of the model during both training and inference.
@@ 5.3.2 Multi-Adapter Fusion on LLMs
We now extend our LLM experiments to the multi-adapter fusion setting. To this end, we create a new setup where we independently train multiple adapters on training sets of individual commonsense reasoning benchmarks, i.e., one adapter each for BoolQ, PIQA, and Arc-Easy. In contrast, each adapter in section 5.3.1 was trained on a combined dataset containing 170K samples from all eight commonsense benchmarks as proposed in [12, 18]. In the present section, the goal is to evaluate how much accuracy drop various adapters experience when we perform multi-adapter fusion. Due to its simplicity towards constructing a mask, we will use SHiRA-WM in the rest of this paper. Further, we explore two settings - overlapping and non-overlapping SHiRA-WM adapters. The overlapping mask consists of top 1% parameters being trained for all tasks. On the other hand, the non-overlapping setting trains the top 1% weights for the first task, next top 1% for the second task and so on. We compare the performance of both LoRA and SHiRA across the multi-adapter fusion of these three tasks.
@@
As shown in Table 4, both overlapping and non-overlapping multi-SHiRA outperforms multi-LoRA on all the three commonsense benchmarks. This is inline with our theoretical analysis in section 4.2 where we suggest that even unstructured sparse SHiRA adapters such as SHiRA-WM would have more orthogonal behavior than LoRA due to high sparsity (see higher AWOR of SHiRA-WM in Fig. 4(right)). In comparison, independently trained LoRA adapters would have no such property and suffer greatly during multi-adapter fusion. As a result, we see that both SHiRA models outperform LoRA by more than 6.5% accuracy on average.
@@
5.4 Generalizing SHiRA to SDXL and DreamBooth
Finally, we extended SHiRA to focus on DreamBooth [23] using a much bigger vision model called SDXL [20].
Figure 6: LoRA- vs. SHiRA-based DreamBooth on SDXL. Prompts for style/subject personalization - left pair: "A picture of a dog in <STYLE:WOODEN-SCULPTURE> style in a bucket", center pair: "A picture of a <SUBJECT:VASE> with flowers", and right pair: "A picture of a sunset in <STYLE:CANVAS> style". Here, "<SUBJECT>" and "<STYLE>" are special identifier tokens for style/subject.
We follow a similar setup as adopted by [2]. Specifically, single-concept datasets and style datasets with five images each collected from the dreambooth dataset [23] and public domains, respectively, were used to train the adapters on one subject – vase – and two style – wooden sculpture and canvas – concepts. For our experiments, we use Stable Diffusion-XL [21] as our base model and train both LoRA and SHiRA adapters with comparable trainable parameters on individual single-concept datasets. During training, special identifier tokens, "<SUBJECT>" or "<STYLE> (e.g., <SBU> as subject token for vase and <SZN> as style token for wooden sculpture and canvas) based prompts are used to finetune subject or style based personalization, respectively. During inference, similar prompts are used to generate images from LoRA- or SHiRA-based DreamBooth.
Fig 6 shows LoRA- and SHiRA-based dreambooth style generated images. Clearly, our proposed adapter produces high quality personalized of target subject in different scenarios. This highlights the broad applicability of our adapter while still preserving the benefits of rapid adapter switching.

### v2

@@ Abstract
Low Rank Adaptation (LoRA) has gained massive attention in the recent generative AI research. One of the main advantages of LoRA is its ability to be fused with pretrained models, adding no overhead during inference. However, from a mobile deployment standpoint, we can either avoid inference overhead in the fused mode but lose the ability to switch adapters rapidly, or suffer significant (up to 30% higher) inference latency while enabling rapid switching in the unfused mode. LoRA also exhibits concept-loss when multiple adapters are used concurrently. In this paper, we propose Sparse High Rank Adapters (SHiRA), a new paradigm which incurs no inference overhead, enables rapid switching, and significantly reduces concept-loss. Specifically, SHiRA can be trained by directly tuning only $1$-$2\%$ of the base model weights while leaving others unchanged. This results in a highly sparse adapter which can be switched directly in the fused mode. We further provide theoretical and empirical insights on how high sparsity in SHiRA can aid multi-adapter fusion by reducing concept loss. Our extensive experiments on LVMs and LLMs demonstrate that finetuning only a small fraction of the parameters in the base model significantly outperforms LoRA while enabling both rapid switching and multi-adapter fusion. Finally, we provide a latency- and memory-efficient SHiRA implementation based on Parameter-Efficient Finetuning (PEFT) Library which trains at nearly the same speed as LoRA while consuming up to $16\%$ lower peak GPU memory, thus making SHiRA easy to adopt for practical use cases. To demonstrate rapid switching benefits during inference, we show that loading SHiRA on a base model can be 5×5$\times$5 ×-16×16$\times$16 × faster than LoRA fusion on a CPU.
@@ 1 Introduction
Low Rank Adaptation (LoRA) [13] is an established technique to tune the behavior of large generative models such as Large Language Models (LLMs) [30, 29] and Stable Diffusion [24, 22]. As the name suggests, LoRA requires very few parameters since it trains low rank projection weights that consume very low memory during the finetuning process while producing excellent results. Moreover, these low rank weights can be fused analytically into the base model, thereby incurring no additional overhead during inference.
@@
Figure 1: Sparse High Rank Adapters (SHiRA): Changing about $1$-$2\%$ weights of the pretrained generative model is often sufficient to achieve high performance. Due to its extreme sparsity, SHiRA enables rapid switching and also reduced concept loss during multi-adapter fusion. In contrast, LoRA modifies majority of parameters when fused, thus prohibiting rapid switching on mobile devices, and also experiences concept loss during multi-adapter fusion. For LoRA, elephant for single “paintings” adapter case has artifacts (extra/broken tusks); bird and knight for multi-adapter case lose “paintings” concept and keep only the “blue fire” effects. SHiRA does not experience these issues.
@@
  • We propose SHiRA, a new high rank adapter paradigm to demonstrate that changing as few as $1$-$2\%$ parameters of the original network is sufficient for adaptation. Our crucial insight is that even the most basic masking criteria (to identify the top $1$-$2\%$ parameters) enable SHiRA to significantly outperform LoRA on diverse vision and language tasks.
@@
  • Finally, on the training side, we provide a PEFT-based latency- and memory-efficient implementation for SHiRA which trains nearly as fast as standard LoRA while consuming $16\%$ lower peak GPU memory. Beyond PEFT, we provide a simple way to turn any trainer into SHiRA finetuning. For inference, we demonstrate that SHiRA weights can be loaded on a CPU up to 5×5$\times$5 ×-16×16$\times$16 × faster than equivalent LoRA fusing, thereby enabling rapid switching.
The rest of this paper is organized as follows: section 2 presents the background and related work. We propose SHiRA in section 3 while describing its theoretical properties in section 4. We then conduct extensive experiments for SHiRA in section 5. Finally, we discuss the key findings in section 6 and conclude the paper in section 7.
@@ 2.2 Related Work
Partial Finetuning. Our work is most closely related to partial finetuning techniques that were mostly proposed in the pre-LoRA era [36, 28, 3, 33, 10]. These methods use a mix of fixed sparse masks [28] or learned masks [36, 10] to finetune a pretrained network. Note that, these techniques have been mostly explored for relatively small language models, and not for recent LLMs and diffusion models. Since the LoRA models exploded in popularity, it has been unclear if other sparse finetuning techniques would achieve comparable results to LoRA on generic adapter tasks, particularly in the vision domain. One significant limitation of partial finetuning, as opposed to LoRA-based methods, is its high GPU memory consumption, making it impractical to be used for large generative models. Consequently, the reduced memory consumption for finetuning was a key factor to LoRA’s success and its widespread adoption. To this end, we provide a memory- and latency-efficient PEFT-based implementation for SHiRA which trains as efficiently as LoRA, thus requiring significantly lower memory consumption compared to prior partial finetuning techniques. Further, we explore the effectiveness of sparse finetuning on both large language and vision models and provide a detailed analysis on rapid switching and multi-adapter fusion of the high rank adapters.
A notable concurrent work is SpIEL [4] which scales partial finetuning to modern LLMs and also has a PEFT implementation that results in comparable speed and memory as LoRA. The main differences between SpIEL and SHiRA are as follows: (i) SpIEL works with dynamic masks while SHiRA uses a static mask. (ii) Dynamic mask in SpIEL requires users to install custom sparse linear layer kernels for the GPUs. In contrast, SHiRA does not require installing any custom kernels and directly works with native Pytorch. Hence, SHiRA’s biggest advantage is its ease of training/inference deployment. (iii) We also analyze multi-adapter fusion properties, e.g., impact of sparsity on orthogonality between adapters, which were not discussed in SpIEL. (iv) Finally, SHiRA demonstrates its effectiveness on both vision and language tasks, whereas SpIEL only discusses the language tasks.
@@
Figure 2: (a) LoRA when fused into the pretrained model modifies all weights and prevents rapid adapter switching. (b) SHiRA does not require additional weights during training but finetunes very few pretrained weights. Our approach relies on a sparse mask for gradient-masking during training. We show that finetuning as low as $1$-$2\%$ parameters is sufficient to achieve high accuracy.
@@ 3.1 Sparse High Rank Adapters (SHiRA)
SHiRA exploits highly sparse trainable parameters in the pretrained model. In its simplest form, our adapter can be trained by masking gradients such that only a fraction of original weights get updated. Specifically, we do not add any new weights to the forward pass like LoRA (see Fig. 2(a)) but rather make a small percentage of existing weights trainable (see Fig. 2(b) top). To this end, we first create an extremely sparse ($\sim$98989898-$99\%$ zeros) mask $\mathcal{M}\in\mathbb{R}^{n\times m}=\{0,1\}^{n\times m}$, where $n,m$ are dimensions of the pretrained weight matrix. $\mathcal{M}$ is then used to mask the gradients during backpropagation using a Hadamard product (see Fig. 2(b) bottom). Thus, very few parameters get updated during training and our adapter consists of just those sparse weights. Concrete gradient masking-based and another latency-/memory-efficient PEFT implementations for SHiRA are discussed in section 3.3.
We consider the following masks $\mathcal{M}$ (only $1$-$2\%$ trainable parameters, see also Appendix B):
@@
    SHiRA-SNIP: The SNIP metric from the pruning literature [19] combines weight magnitude and gradient strategies, i.e., SNIP equals magnitude of the gradient times the weight.
@@ 3.2 Rapid Adapter Switching, Multi-Adapter Fusion, and High Rank
Since very few base weights change during the SHiRA training, we can simply extract them out and store them as sparse weights and their indices (see Fig. 3(a)). Hence, SHiRA is comparable to LoRA in model size but overwrites only a fraction of the pretrained weights at inference time. In contrast, LoRA fuses into base weights as $W_{new}=W+AB$ and changes the entire weight. Note that, we do not actually need to fuse SHiRA but rather just need to overwrite the modified value at the correct index in the pretrained weight tensor. This enables rapid switching on resource-constrained devices. To verify that SHiRA indeed provides rapid switching benefits compared to LoRA, we provide an optimized implementation based on scatter_op to overwrite base model weights instead of fusing them like LoRA. We demonstrate that on a CPU, weight loading for SHiRA adapters can be up to $\bm{5\times}$-$\bm{16\times}$ faster than equivalent LoRA fusing for inference (see Appendix C and Fig 7).
Next, we discuss multi-adapter fusion in SHiRA. Given two adapters $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ with sparse masks $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, we ask the following questions: (i) What is the impact of sparsity on relative interference between adapters in the multi-adapter setting? (ii) Is it possible to create masks that result in nearly orthogonal SHiRA weights so they do not significantly interfere with each other at inference time?
@@
Finally, since we do not have any low rank weights in the forward pass, our proposed adapters can be high rank albeit highly sparse. We theoretically analyze the rank vs. sparsity properties in section 4.
@@ 4 Theoretical Insights for SHiRA
4.1 Rank vs. Sparsity
@@ 4.2 Adapter Weight Orthogonality in Multi-Adapter Fusion
Non-overlapping SHiRA-Struct adapters are nearly orthogonal: AWOR for non-overlapping SHiRA-Struct adapters is at most the sum of sparsity of individual adapters. Since all SHiRA masks are highly sparse, $\mathcal{A}_{1}^{T}\mathcal{A}_{2}$ has a lot of zeros, thus making the adapters nearly orthogonal.
@@
Remark 1. The orthogonality property shown here can lead to disentangled representation for adapter outputs before they merge into the base model. However, this property does not hold for other SHiRA masks that do not have a regular sparsity pattern like SHiRA-Struct even if other SHiRA strategies are still more orthogonal than LoRA weights (e.g., see SHiRA-WM AWOR in Fig. 4(right)). Interestingly, for unstructured sparse masks like SHiRA-WM, SHiRA-Grad, SHiRA-SNIP, etc., both overlapping and non-overlapping adapters have similar orthogonality properties. We discuss this in more detail in section 5.3.2. Finally, this analysis only focuses on orthogonality of adapter weights and not on orthogonality of subspaces. We leave the subspace analysis of SHiRA for future work.
@@ 5.2.1 Impact of Various SHiRA Masks
Next, we compare various types of SHiRA masks in Fig. 5. Clearly, all SHiRA schemes produce impressive images for different prompts and significantly outperform LoRA. We further quantify the image quality using HPSv2 for each of the masks. The results are presented in Table 1. As evident, all variants of SHiRA consistently achieve superior or similar HPSv2 scores than LoRA, especially for larger $\alpha$ (see details on scaling factor $\alpha$ in Appendix I). More results are provided in Appendices J and K: see Table 10 and Fig.  10,  1112.
@@ 5.2.2 SHiRA Adapters aid Multi-Adapter Fusion
We now validate the effectiveness of various SHiRA schemes on multi-adapter fusion. The right two columns in Fig. 1 and Fig. 5 show our results. SHiRA is clearly better at capturing both concepts than LoRA. For example, both bird and knight images in Fig. 1 generated with LoRA lose most of the paintings concept. Similarly, for the fox image in Fig. 5, LoRA does not show significant bluefire concept. In contrast, SHiRA-Struct and SHiRA-SNIP consistently perform well on many different prompts and produce exceptional images for multi-adapter fusion. Please refer to Appendix K.1 (Fig. 101112, and  13) for additional results. For certain classes that were not included in the training set for both adapters (e.g., see Koala in Fig. 1012, and 13 in Appendix), we observe that LoRA produces significant artifacts whereas SHiRA generates high quality images.
@@ 5.3.1 Single Adapter SHiRA Finetuning
Similar to vision results, we demonstrate the effectiveness of SHiRA on language tasks. For our experiments, each adapter (i.e., weight-magnitude, gradient-magnitude, and SNIP based SHiRA) is trained on the combined 170K sample commonsense reasoning dataset released by [14, 20]. Similar to [20], we train our SHiRA adapters for 3 epochs and compare it against the LoRA baselines. As shown in Table 2, various SHiRA adapters outperform LoRA by 1.9-2.7% on an average on LLaMA-7B. Importantly, SHiRA only modifies 1% base parameter weights as compared to 66.72% (4.5B weights) changed by LoRA in the fused mode, thus enabling rapid switching on edge devices. Interestingly, we found that SHiRA-Struct does not perform well on language tasks likely because it is a rank 1 + diagonal adapter and may not have sufficient expressive power.
@@
Table 2: Evaluation of LLaMA-7B on Commonsense Reasoning. WG and HS denote WinoGrande and HellaSwag, respectively. %C represents parameters changed in the fused mode. ($\uparrow$): the higher the better. Green denotes improvement. Trained by masking a high-rank DoRA with a WM mask of top 1% weights, thus changing only 1% of the model during both training and inference.
@@ 5.3.2 Multi-Adapter Fusion on LLMs
We now extend our LLM experiments to the multi-adapter fusion setting. To this end, we create a new setup where we independently train multiple adapters on training sets of individual commonsense reasoning benchmarks, i.e., one adapter each for BoolQ, PIQA, and Arc-Easy. In contrast, each adapter in section 5.3.1 was trained on a combined dataset containing 170K samples from all eight commonsense benchmarks as proposed in [14, 20]. In the present section, the goal is to evaluate how much accuracy drop various adapters experience when we perform multi-adapter fusion. Due to its simplicity towards constructing a mask, we will use SHiRA-WM in the rest of this paper. Further, we explore two settings - overlapping and non-overlapping SHiRA-WM adapters. The overlapping mask consists of top 1% parameters being trained for all tasks. On the other hand, the non-overlapping setting trains the top 1% weights for the first task, next top 1% for the second task, and so on. We compare the performance of both LoRA and SHiRA across the multi-adapter fusion of these three tasks.
@@
As shown in Table 4, both overlapping and non-overlapping multi-SHiRA outperform multi-LoRA on all three commonsense benchmarks. This is inline with our theoretical analysis in section 4.2 where we suggest that even unstructured sparse SHiRA adapters such as SHiRA-WM would have more orthogonal behavior than LoRA due to high sparsity (see higher AWOR of SHiRA-WM in Fig. 4(right)). In comparison, independently trained LoRA adapters would have no such property and suffer greatly during multi-adapter fusion. As a result, we see that both SHiRA models outperform LoRA by more than 6.5% accuracy on average. Further analysis of the properties of these trained adapters is discussed in Appendix K.3 (see Table 13 and Fig. 9).
@@
5.4 Content/Style Personalization: Generalizing SHiRA to SDXL and DreamBooth
Finally, we extend SHiRA to focus on DreamBooth [25] using a much bigger vision model called SDXL [22].
Figure 6: LoRA- vs. SHiRA-based DreamBooth on SDXL. Prompts for content/style personalization - left pair: "A picture of a dog in <STYLE:WOODEN-SCULPTURE> style in a bucket", center pair: "A picture of a <CONTENT:VASE> with flowers", and right pair: "A picture of a sunset in <STYLE:CANVAS> style". Here, "<CONTENT>" and "<STYLE>" are special identifier tokens for content/style.
We follow a similar setup as adopted by [2]. Specifically, one content (vase) and two style (wooden sculpture and canvas) datasets with five images each were collected from the DreamBooth dataset [25] and public domains, respectively. These datasets were used to train various content and style adapters. For our experiments, we use SDXL [23] as our base model and train both LoRA and SHiRA adapters with comparable trainable parameters on individual single-concept datasets. During training, prompts containing special identifier tokens like "<CONTENT>" or "<STYLE>" (e.g., <SBU> as content token for vase and <SZN> as style token for wooden sculpture and canvas) are used to finetune the SDXL network for content or style personalization, respectively. During inference, similar prompts are used to generate images from LoRA- or SHiRA-based DreamBooth.
Fig 6 shows DreamBooth generated images for LoRA and SHiRA. Clearly, our proposed adapter produces high quality personalized images of target concept in different scenarios. This highlights the broad applicability of our adapter while still preserving the benefits of rapid adapter switching.
6 Discussion
To summarize our main contributions, we highlight that SHiRA – when used with even the most basic pruning metrics (such as weight- or gradient-magnitude, SNIP, structured masks, etc.) – significantly outperforms LoRA on a variety of large-scale tasks in both large vision and large language domains. For LVM style transfer applications, we found that SHiRA-Struct is the most effective masking technique due to its special orthogonality properties that aid multi-adapter fusion. However, SHiRA-SNIP and SHiRA-Grad are not too far behind and achieve competitive performance as SHiRA-Struct. On the LLM commonsense reasoning side, SHiRA-SNIP is the best strategy out of the masking techniques we have considered in this work. Specifically, SHiRA-Struct did not achieve good results on the more complex commonsense reasoning tasks since it is a combination of a rank-1 + a highly sparse diagonal adapter. SHiRA-Grad on LLMs is about 0.8% worse accuracy than SHiRA-SNIP (76.6% vs. 77.4% average accuracy on commonsense reasoning for LLaMA-1). Therefore, in conclusion, for the applications/fields and the masking techniques considered in this paper, SHiRA-SNIP works well across both language and vision domains. Hence, we recommend that SHiRA-SNIP is one of the strongest candidates that we have considered for sparse finetuning.